27 January 2014 – Monday

Backups:

  • Attached and configured serial cables for root login on stornode3 and stornode4.  Attached a serial cable to bujbod1, but I haven’t found a way to let root login without making it the only console.
  • 20TB left on the ZFS pool “athens” that I’ve been working to clear off for over a month.  Might actually get it done this week.  Trying to get the weekend’s full save processes to complete – have bailed out the staging from “athens” to free up a tape drive.
  • Windows Fileserver, COM/IS MED9300, and Structured Biology weekend full groups are still running this morning.
  • Because of messages that stages are waiting from the scripted stage process, I have doubled the max number of nsrmmd’s (from 4 to 8) that can run for the afflicted NFS mounts on stornode2, stornode3, and stornode4 (everything except the PeopleSoft volume).

LDAP:

  • One update issue from over the weekend.  Two former students showed up in Friday night’s update.  Dancing fool mode, engaged.
  • One email alias flopped around per request from Account Services.

Tags:

Leave a Reply

You must be logged in to post a comment.